how do you find the y-intercept to an equation like this y=3/5x-4





Answer by stanbon(75887)

find the y-intercept to an equation like this y=3/5x-4
-----------------
Let x=0 and solve for y.
y=(3/5)(0)-4
y=0-4
y=-4
The y-intercept is -4
Cheers,
stan H.